OpenStudy (anonymous):

PLEASE HELP! QUICK QUESTION! 2x+ay=4 ax+8y=8 What value(s) of "a" will give no solution and infinite solution?

OpenStudy (anonymous):

is the question correct?

OpenStudy (anonymous):

are those supposed to be two equations describing the same variable?! if so, no value of a would satisfy both equations

OpenStudy (anonymous):

my bad, the equation is 2x+ay=4 and ax+8y=8

OpenStudy (anonymous):

first, combine equations to make it in the form ax^2 + bx + c = 0 (quadratic form). where 'x' is 'a' in this case

OpenStudy (anonymous):

then, you can set up conditions to check when the solution will not exist or be infinite

OpenStudy (triciaal):

another approach to have infinite solutions equation 1 would be equal to equation 2 to make them equal multiply the first equation by 4 8 = 8 so 4x + 2ay = 8 and ax + 8y = 8 therefore when a = 4

OpenStudy (anonymous):

what about no solutions?

OpenStudy (triciaal):

what do you think?

OpenStudy (anonymous):

i'd say when a -0

OpenStudy (anonymous):

a=0

OpenStudy (triciaal):

I did not post because not sure. The no solution situation is when the solution for equation one when substituted in equation 2 will not make the equation true.

OpenStudy (triciaal):